# How Generalizable Is My Behavior Cloning Policy? A Statistical Approach to Trustworthy Performance Evaluation

### v1

@@ V Confidence Bounds - Continuous Metric
where $\mathbf{1}(x)$ evaluates to $1$ if $x$ is true and 00 otherwise. A one-sided Kolmogorov-Smirnov (KS) statistic is
@@
If $F(x)$ is continuous, then the distribution of $D_{n}^{-}$ is
@@
Then, given i.i.d. samples $X_{1:n}$ from some $F(x)$, we have the following upper confidence bound on the CDF,
@@
The tightness of this bound is measured by $\epsilon^{*}$, the offset from the empirical CDF. The $\epsilon^{*}$ chosen by Eq. 18c is optimal, if made any smaller then Eq. 18a would not hold. Practical computation of $\epsilon^{*}$ given $n$ and $\alpha$ is easily done via a bisection search. Given $\alpha$ and $\epsilon^{*}$, it is then simple to compute the minimum number of policy rollouts $n$ needed to meet these specifications. One can also use the DKW inequality [10, 11] to obtain a similar bound, which takes the same form as Eq. 18b except with $\epsilon^{*}~{}=~{}\sqrt{-\ln{\alpha}/(2n)}$. However, the DKW bound is always less tight as shown in Fig. 4.
@@ VI Experiments
We demonstrate our approach to evaluating BC policies in both simulation and real-world robotic manipulation settings. The performance of BC policies in OOD settings is hard to predict since it is difficult to characterize what aspects of the policy are invariant or sensitive to particular distribution shifts. This motivates the need for rigorous statistical evaluation. Thus, all our experiments evaluate policies in OOD settings. In simulation, we run many evaluations and show the empirical confidence and tightness of the bounds agree with the theory. In hardware, we (i) investigate the degree to which a learned policy generalizes to OOD settings, and (ii) compare the OOD generalization of two learned policies. Our approach requires i.i.d. performance scores; a sufficient (but not necessary) condition to ensure this is for the initial observations to be i.i.d. from some (potentially unknown) distribution. In simulation, this can be trivially satisfied. In hardware, we took efforts to minimize the impact of any time-varying conditions that may undermine this assumption. Because our bounds are strictly tighter than those of Clopper-Pearson and DKW, we report the results for our bounds only.
@@ VI-A Simulation Results
We consider five diffusion policies from [1] for visuomotor robot manipulation that have shown recent SOTA performance, each trained for a separate task. Fig. 5 shows snapshots of each task in the training environments.
After training, we modified each environment as follows,
@@
From the figure we see that the distribution shifts for the Can and Square environments were quite harmful to the performance of the policy, while those for the other tasks were benign. Our bounds tend to accurately reflect these trends. These results reflect the unintuitive nature about when learned policies generalize. For instance, it is not clear why changing the color of the manipulated object is harmful in some cases (Can and Square) but benign in other cases (Lift and Transport). Lastly, the Monte Carlo estimates of the reward distributions are optimistic about half the time, making them unfit to be interpreted as bounds. Table I shows good agreement between the empirical confidence and tightness estimates with their theoretical values.
@@ VI-B Generalization of a Single Policy
The bounds we find support our hypotheses, and importantly also inform us as to the degree of performance we can assert at a $95\%$ confidence level. If a user requires a $60\%$ success rate, then we could assert with $95\%$ confidence that the benign setting is suitable for deployment whereas we could not for the harmful setting. We may then choose to retrain the policy with data from this setting and re-evaluate.
@@ VI-C Comparing the Generalization of Two Policies
In this experiment we compare the performance of two learned visuomotor manipulation policies. Specifically, we apply our statistical bounds to the recent results from [43], where the authors compare their RT-2 policy to a VC-1 policy in three settings designed to test emergent capabilities in symbol understanding, reasoning, and human recognition. We compute a lower confidence bound on the success rate of the RT-2 policy ($\underline{p}_{\text{RT-2}}$) and an upper confidence bound on the success rate of the VC-1 policy ($\overline{p}_{\text{VC-1}}$). Note, $\overline{p}_{\text{VC-1}}=1-\underline{q}_{\text{VC-1}}$, where $\underline{q}_{\text{VC-1}}$ is a lower confidence bound on the failure rate. We construct the bounds with confidence level $0.975$, to ensure a joint confidence level of $0.95$,
@@
In Fig. 8, we show the bounds for each policy in each evaluation setting. We find in each case there is enough evidence to conclude (at a $95\%$ confidence level) that the RT-2 policy outperforms the VC-1 policy. For some settings the gap between the bounds is larger, this may be due to the actual success rates being further apart, or an effect of sample size. With separate bounds for each policy, we can also conclude how well each policy performs individually, an important aspect that Fisher’s exact test does not provide [36]. Finally, this approach easily extends to the case of continuous rewards. A lower bound on the CDF is computed similarly to an upper bound by noting that $D_{n}^{+}=\sup_{x}\ F_{n}(x)-F(x)$ has the same distribution as $D_{n}^{-}$ (Eq. 16[41].
@@ VII Conclusion
We build upon classical statistical techniques to rigorously evaluate BC policies. We use worst-case confidence bounds to measure policies’ OOD generalization of to unseen environments. For binary and continuous performance metrics, we place confidence bounds on the entire distribution (i.e. CDF) of a policy’s performance. These bounds hold with a user-specified confidence level, tightness, and are constructed from as few policy rollouts as necessary. For binary performance metrics, we introduce a solution for tractably evaluating the tightness via mixed-monotonic programming. Finally, in three experiments we show both the validity of our approach as well as how the approach can be used to make judgements about policy generalization in OOD settings.
In this paper, the number of policy rollouts is determined by a budget or by constraints on confidence and tightness. However, our ideas can be extended to use confidence sequences where the number of rollouts is not determined beforehand [44]. It may save on testing costs to stop running policy rollouts when enough evidence has been collected to conclude that the policy surpasses some performance threshold. Another direction is to evaluate performance during a single policy rollout, rather than multiple i.i.d. rollouts. As a final remark, the utility of any statistical analysis depends on the researchers being forthright in their data collection, assumptions/modeling, and interpretation of significance [45].

### v2

@@ V Confidence Bounds - Continuous Metric
where $\mathbf{1}(x)$ evaluates to $1$ if $x$ is true and 00 otherwise. A one-sided Kolmogorov-Smirnov (KS) statistic [36] is
@@
By [42], if $F(x)$ continuous, then the distribution of $D_{n}^{-}$ is
@@
Then, given i.i.d. samples $X_{1:n}$ from some $F(x)$, we have the following upper confidence bound on the CDF [42],
@@
The tightness of this bound is measured by $\epsilon^{*}$, the offset from the empirical CDF. The $\epsilon^{*}$ chosen by Eq. 18c is optimal, if made any smaller then Eq. 18a would not hold. $\epsilon^{*}$ is easily computed via a bisection search. Since $\epsilon^{*}$ is decreasing in $n$, to determine $n$ given a desired $\alpha$ and $\epsilon^{*}$, one fixes $\alpha$ and computes the smallest $n$ (via a bisection search) which yields an $\epsilon^{*}$ below the desired value. One can also use the DKW inequality [10, 11] to obtain a similar bound, except with $\epsilon^{*}~{}=~{}\sqrt{-\ln{\alpha}/(2n)}$. However, the DKW bound is always less tight as shown in Fig. 4.
@@ VI Experiments
We demonstrate our approach to evaluating BC policies in both simulation and real-world robotic manipulation settings. Performance of BC policies in OOD settings is hard to predict since it is difficult to characterize what aspects of the policy are invariant or sensitive to particular distribution shifts. This motivates the need for rigorous statistical evaluation. Thus, all of our experiments evaluate policies in OOD settings. In the absence of assumptions on the nature of the distribution shift, it is necessary to directly test the policies in the OOD settings.
In simulation, we run many evaluations and show the empirical confidence and tightness of the bounds agree with the theory. In hardware, we (i) investigate the degree to which a learned policy generalizes to OOD settings, and (ii) compare the OOD generalization of two learned policies. Our approach requires i.i.d. performance scores; a sufficient condition to ensure this is for the initial observations to be i.i.d. from some (potentially unknown) distribution. In simulation, this is easily satisfied. In hardware, we took efforts to eliminate any time-varying conditions that may undermine this assumption. Because our bounds are strictly tighter than those of Clopper-Pearson and DKW, we report the results for our bounds only.
@@ VI-A Simulation Results
We consider five diffusion policies from [1] for visuomotor manipulation that have shown recent SOTA performance, each trained for a separate task. Fig. 5 shows training environments. After training, we modified each environment as follows,
@@
From the figure we see that the distribution shifts for the Can and Square environments were quite harmful to the policy performance, while those for the other tasks were benign. Our bounds tend to accurately reflect these trends. These results reflect the unintuitive nature about when learned policies generalize, e.g., it is not clear why changing the color of the manipulated object is harmful in some cases (Can and Square) but benign in other cases (Lift and Transport). Lastly, the Monte Carlo estimates of the distributions are overly optimistic, making them unfit to be interpreted as bounds. Table I shows good agreement between the empirical confidence and tightness estimates with their theoretical values.
@@ VI-B Generalization of a Single Policy
The bounds we find support our hypotheses, and also inform us as to the degree of performance we can assert at a $95\%$ confidence level. If a user requires a $60\%$ success rate, then we could claim with $95\%$ confidence that the benign setting is suitable for deployment whereas we could not for the harmful setting. We may then choose to retrain the policy with data from this setting and re-evaluate.
@@ VI-C Comparing the Generalization of Two Policies
In this experiment we compare two learned visuomotor manipulation policies from [44], where the authors test their RT-2 policy against a VC-1 policy in three settings designed to test emergent capabilities in symbol understanding, reasoning, and human recognition. We compute a lower confidence bound on the success rate of the RT-2 policy ($\underline{p}_{\text{RT-2}}$) and an upper confidence bound on the success rate of the VC-1 policy ($\overline{p}_{\text{VC-1}}$). Note, $\overline{p}_{\text{VC-1}}=1-\underline{q}_{\text{VC-1}}$, where $\underline{q}_{\text{VC-1}}$ is a lower confidence bound on the failure rate. Each bound has confidence level $0.975$, to ensure a joint confidence of $0.95$,
@@
In Fig. 8, we show the bounds for each policy in each evaluation setting. We find in each case there is enough evidence to conclude (at a $95\%$ confidence level) that the RT-2 policy outperforms the VC-1 policy. For some settings the gap between the bounds is larger, this may be due to the actual success rates being further apart, or an effect of sample size. With separate bounds for each policy, we can also conclude how well each policy performs individually, an important aspect that standard A/B testing approaches (like Fisher’s exact test) do not provide [35]. Finally, this approach easily extends to the case of continuous rewards. A lower bound on the CDF is easily computed since $D_{n}^{+}=\sup_{x}\ F_{n}(x)-F(x)$ has the same distribution as $D_{n}^{-}$ (Eq. 16[42].
@@ VII Conclusion
The primary contribution of this paper is utilizing uncommon but optimal statistical bounds to efficiently measure policies’ OOD generalization to unseen environments from few samples. For binary and continuous performance metrics, we place confidence bounds on the entire distribution (i.e. CDF) of a policy’s performance. To this end, a secondary contribution of this paper is the computation of MES for the UMA and Clopper-Pearson lower confidence bounds, which has applicability beyond robotics. The final contribution of our paper is an open-source implementation of the UMA lower confidence bound and associated MES computation, both of which were absent from existing statistical software. Finally, in three experiments we show both the validity of our approach as well as how the approach can be used to make judgements about policy generalization in OOD settings. Although we focus our discussion and experiments on BC policies, our approach can be applied to any black-box policy.
The bounds presented in this paper are valid for the environment the samples were drawn from. One can extend the analysis in [26] to understand the sensitivity of the bounds’ confidence level to distribution shifts, or to construct bounds which are robust to prescribed levels of distribution shift. In our paper, the number of policy rollouts is determined by a budget or by constraints on confidence and tightness. However, our ideas can be extended to use confidence sequences where the number of rollouts is not determined beforehand [37]. It may save on testing costs to stop when enough evidence has been collected to conclude that the policy surpasses some performance threshold. Another direction is to evaluate performance during a single policy rollout, rather than multiple i.i.d. rollouts. As a final remark, when performing statistical evaluations, researchers should be forthright in their data collection, assumptions/modeling, and interpretation of results.